OpenStudy (anonymous):

Ana bought two melons and one pineapple for $8.96. If she had purchased one melon and two pineapples, she would have spent $1.49 more. What is the price of a melon? a pineapple?

OpenStudy (mertsj):

2m+p=8.96 m+2p=10.45 m=10.45-2p 2(10.45-2p)+p=8.96 20.90-4p+p=8.96 -3p=-11.94 p=3.98 m=10.45-2(3.98) m=2.49

OpenStudy (anonymous):

Hello again Mertsj

OpenStudy (mertsj):

Hi

OpenStudy (anonymous):

you are very helpful now is that answer laid out one after another correct it isn't all bunched up